# Score Distillation of Flow Matching Models

### v1

@@ 5 Conclusion
In this work, we revisited the theoretical foundations of diffusion and flow matching models, showing that under Gaussian assumptions, their optimal solutions are equivalent despite differences in loss weighting and practical implementations. Building on this unified perspective, we demonstrated that score distillation—originally developed for diffusion models—can be effectively and robustly extended to flow matching models without requiring model-specific adaptations or teacher finetuning. Through the use of few-step Score identity Distillation (SiD), we successfully distilled a wide range of pretrained text-to-image flow matching models, including SANA, SD3, SD3.5, and FLUX.1-dev, into efficient four-step generators. Our approach uses a single, shared codebase and training configuration across models of varying architectures and parameter scales, showcasing the generality and stability of score distillation in this new context. These findings not only clarify misconceptions in prior work regarding the applicability of score distillation to flow-based models, but also open new directions for compressing, accelerating, and enhancing modern text-to-image generators. By bridging the gap between diffusion and flow matching, our work provides a solid theoretical and empirical foundation for future research on unified generative modeling and fast sampling strategies.
@@
We will release the codebase and training scripts to facilitate reproduction of all experiments. All algorithmic derivations are detailed in the main text, while hyperparameter settings and precision configurations (AMP vs. BF16) are reported in Tables 3 and 4 of the Appendix.
@@ Appendix A Use of Large Language Models (LLMs)
Large Language Models (LLMs) were used to improve grammar, clarity, and readability of the text. They also assisted with code debugging and annotation.
@@ Appendix B Related Work
Flow matching has recently emerged as a promising alternative for generative modeling (Liu et al., 2022b; Lipman et al., 2022; Albergo et al., 2023). A key example is rectified flow (Liu et al., 2022b), also known as flow matching with an optimal transport path (Lipman et al., 2022). Rectified flow encourages straighter trajectories between noise and data, reducing the number of function evaluations (NFEs) needed for sampling and enabling one- or few-step generation via ReFlow (Liu et al., 2022b). Another representative approach is TrigFlow (Lu & Song, 2025), now the preferred framework for continuous consistency distillation and successfully applied by Chen et al. (2025) to develop SANA-Sprint, which distills SANA T2I models after finetuning rectified flow teachers into TrigFlow. In contrast, our method works directly with SANA models trained under either rectified flow or TrigFlow, without requiring such finetuning.
Although originally proposed as a faster and simpler alternative to diffusion, recent theoretical insights have shown that, under Gaussian assumptions, rectified flow is fundamentally equivalent to diffusion: training a Gaussian noise-based rectified flow model is mathematically equivalent to training a Gaussian diffusion model, and their corresponding SDE/ODE sampling procedures are interchangeable (Albergo et al., 2023; Kingma & Gao, 2023; Ma et al., 2024; Gao et al., 2024; Geffner et al., 2025), and thus the distillation techniques proposed for diffusion models can be adapted to Gaussian-based rectified flow, such as consistency models (Yang et al., 2024; Lu & Song, 2025). Nevertheless, practical differences remain, such as in noise schedules, loss formulations, and network architectures.
@@ Appendix C Weight Normalized Time Schedule
Comparing the velocity-prediction loss shown in (9) and the $\epsilon$-prediction loss shown in (7), we observe that they differ by a time-dependent scaling factor $\alpha_{t}^{2}$. However, as discussed earlier, one must consider both the distribution of $p(t)$ and the weighting function $w_{t}$ when evaluating how each $t$ contributes to the overall loss. From this perspective, while the DDPM schedule appears to place more emphasis on values of $t$ closer to one (i.e., by sampling them more frequently), it down-weights the corresponding $x_{0}$-prediction loss more than the SANA schedule does.
Appendix D Hyperparameter Settings
@@
Appendix E Additional Qualitative Examples
Figure 4: Qualitative results of the four-step SiD-DiT generator distilled from FLUX-1.DEV in a data-free setting ($i.e.$, without access to additional training data).
@@ Appendix E Additional Qualitative Examples
Appendix F Prompt Details

### v2

@@ 5 Conclusion
In this work, we revisited the theoretical foundations of diffusion and flow matching models, showing that under Gaussian assumptions, their optimal solutions are equivalent despite differences in loss weighting and practical implementations. Building on this unified perspective, we demonstrated that score distillation—originally developed for diffusion models—can be effectively and robustly extended to flow matching models without requiring model-specific adaptations or teacher finetuning. Through the use of few-step Score identity Distillation (SiD), we successfully distilled a wide range of pretrained text-to-image flow matching models, including SANA, SD3, SD3.5, and FLUX.1-dev, into efficient four-step generators. Our approach uses a single, shared codebase and training configuration across models of varying architectures and parameter scales, showcasing the generality and stability of score distillation in this new context. These findings not only clarify misconceptions in prior work regarding the applicability of score distillation to flow-based models, but also open new directions for compressing and accelerating modern text-to-image generators. By bridging the gap between diffusion and flow matching, our work provides a solid theoretical and empirical foundation for future research on unified generative modeling and fast sampling strategies.
@@
To facilitate reproduction of all experiments, we release the full codebase and training scripts on our project page: https://yigu1008.github.io/SiD-DiT. All algorithmic derivations are detailed in the main text, while hyperparameter settings and precision configurations (AMP vs. BF16) are reported in Tables 4 and 5 of the Appendix.
@@ Appendix A Use of Large Language Models (LLMs)
Large Language Models (LLMs) were used to improve grammar, clarity, and readability of the text. They also assisted with code debugging, annotation, and anonymization.
@@ Appendix B Related Work
Flow matching has recently emerged as a promising alternative for generative modeling (Liu et al., 2022b; Lipman et al., 2022; Albergo et al., 2023). A key example is rectified flow (Liu et al., 2022b), also known as flow matching with an optimal transport path (Lipman et al., 2022). Rectified flow encourages straighter trajectories between noise and data, reducing the number of function evaluations (NFEs) needed for sampling and enabling one- or few-step generation via ReFlow (Liu et al., 2022b). Another representative approach is TrigFlow (Lu & Song, 2024), now the preferred framework for continuous consistency distillation and successfully applied by Chen et al. (2025) to develop SANA-Sprint, which distills SANA T2I models after finetuning rectified flow teachers into TrigFlow. In contrast, our method works directly with SANA models trained under either rectified flow or TrigFlow, without requiring such finetuning.
Although originally proposed as a faster and simpler alternative to diffusion, recent theoretical insights have shown that, under Gaussian assumptions, rectified flow is fundamentally equivalent to diffusion: training a Gaussian noise-based rectified flow model is mathematically equivalent to training a Gaussian diffusion model, and their corresponding SDE/ODE sampling procedures are interchangeable (Albergo et al., 2023; Kingma & Gao, 2023; Ma et al., 2024; Gao et al., 2024; Geffner et al., 2025), and thus the distillation techniques proposed for diffusion models can be adapted to Gaussian-based rectified flow, such as consistency models (Yang et al., 2024; Lu & Song, 2024). Nevertheless, practical differences remain, such as in noise schedules, loss formulations, and network architectures.
@@ Appendix C Weight Normalized Time Schedule
Comparing the $v$-prediction loss shown in (9) and the $\epsilon$-prediction loss shown in (7), we observe that they differ by a time-dependent scaling factor $\alpha_{t}^{2}$. However, as discussed earlier, one must consider both the distribution of $p(t)$ and the weighting function $w(t)$ when evaluating how each $t$ contributes to the overall loss. From this perspective, while the DDPM schedule appears to place more emphasis on values of $t$ closer to one (i.e., by sampling them more frequently), it down-weights the corresponding $x_{0}$-prediction loss more than the SANA schedule does.
Appendix D Algorithmic Pseudo-Code
Algorithm 1 Score Distillation of DiT-Based Flow-Matching T2I Generation
1:Input: Pretrained DiT $v_{\phi}$, generator DiT $G_{\theta}$, fake score DiT $v_{\psi}$, $t_{\text{init}}=999$, training timestep distribution $p(t)=\operatorname{Logit}{\mathcal{N}}\bigl(t;\mu,\sigma\bigr)$, learning rate $\eta$.
2:Initialization: $\theta\leftarrow\phi,\ \psi\leftarrow\phi$
3:repeat
4:  Update Fake Score
5:  Sample $\bm{z}\sim\mathcal{N}(0,{\bf I})$ and set $\bm{x}_{g}\leftarrow G_{\theta}(t_{\text{init}},\bm{z})$
6:  Sample $t\sim p(t)$ and $\bm{\epsilon}_{t}\sim\mathcal{N}(0,{\bf I})$, and set $\bm{x}_{t}\leftarrow(1-t)\bm{x}_{g}+t\bm{\epsilon}_{t}$
7:  Use (15) to compute CFG-modified $f_{\psi}(\bm{x}_{t},{\bm{c}})$ based on flow prediction $v_{\psi}(\bm{x}_{t},{\bm{c}})$
8:  Update $\psi$ with:
9:  $\mathcal{L}_{\psi}=\bigl\|f_{\psi}(\bm{x}_{t},{\bm{c}})-\bm{x}_{g}\bigr\|_{2}^{2}$,  $\psi\leftarrow\psi-\eta\nabla_{\psi}\mathcal{L}_{\psi}$
10:  Update Generator
11:  Sample $t\sim p(t)$ and compute $\omega_{t}$ using any combination listed in the caption of Figure 2.
12:  Unless specified otherwise, we use $p(t)=\operatorname{Logit}\mathcal{N}(\ln 2,1.6^{2})$ and $w_{t}=1-t$ for all models in the paper.
13:  Sample generator update step uniformly at random from $k\in\{1,2,3,4\}$.
14:  Generate ${\bm{x}}_{g}^{(k)}$ as in (16) and forward diffuse ${\bm{x}}_{t}^{(k)}$ as in (17).
15:  Compute $f_{\phi}({\bm{x}}_{t}^{(k)},{\bm{c}})$ based on flow prediction $v_{\phi}({\bm{x}}_{t}^{(k)},{\bm{c}})$ using (15).
16:  Compute $f_{\psi}({\bm{x}}_{t}^{(k)},{\bm{c}})$ based on flow prediction $v_{\psi}({\bm{x}}_{t}^{(k)},{\bm{c}})$ using (15).
17:  Update $G_{\theta}$ with:
18:  $\mathcal{L}_{\theta}({\bm{x}}_{t}^{(k)})=w_{t}\big(f_{\phi}({\bm{x}}_{t}^{(k)},t_{k},{\bm{c}})-f_{\psi}({\bm{x}}_{t}^{(k)},t_{k},{\bm{c}})\big)^{\!\top}\big(f_{\psi}({\bm{x}}_{t}^{(k)},t_{k},{\bm{c}})-{\bm{x}}_{g}^{(k)}\big)$
19:  $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}_{\theta}$
20:until the FID plateaus or the training budget is exhausted
21:Output: $G_{\theta}$
Appendix E Detailed GenEval Scores
[…]
Model GenEval single_object counting color_attr colors position two_object
SANA 0.6B
SANA (Xie et al., 2025) 0.64087 1.0000 0.6281 0.4225 0.8910 0.2044 0.6992
SiD-DiT (SANA) 0.65212 0.9812 0.5969 0.4475 0.8484 0.2800 0.7727
[…]
SANA 1.6B
SANA (Xie et al., 2025) 0.65501 0.9969 0.5782 0.3925 0.8856 0.2770 0.7998
SiD-DiT (SANA) 0.67013 0.9781 0.5437 0.4950 0.8590 0.3000 0.8333
[…]
SANA TrigFlow 0.6B
SANA Sprint Teacher 0.78018 1.0000 0.7000 0.5575 0.9069 0.5975 0.9192
SANA Sprint (TrigFlow, 1 step) 0.77074 0.9938 0.6469 0.4650 0.8883 0.5300 0.8005
SANA Sprint (TrigFlow, 4 steps) 0.76591 1.0000 0.6812 0.5150 0.8803 0.6300 0.8889
SiD-DiT (SANA, TrigFlow) 0.76289 1.0000 0.5594 0.5050 0.8936 0.5700 0.9116
[…]
SANA TrigFlow 1.6B
SANA Sprint Teacher 0.77571 1.0000 0.6719 0.5725 0.8830 0.5875 0.9394
SANA Sprint (TrigFlow, 1 step) 0.76796 0.9938 0.5219 0.5025 0.8936 0.5425 0.8535
SANA Sprint (TrigFlow, 4 steps) 0.76769 1.0000 0.5844 0.5275 0.9149 0.5525 0.8889
SiD-DiT (SANA, TrigFlow) 0.77421 0.9875 0.6219 0.5350 0.9096 0.6350 0.9369
SiD2-DiT (SANA, TrigFlow) 0.76829 0.9906 0.6562 0.5175 0.8617 0.5775 0.9015
Appendix F Hyperparameter Settings
@@
Table 6: Estimated training cost of SiD-DiT with different teacher models, measured in thousands of images processed (k imgs) and in estimated machine hours, shown for both training to the final checkpoint and for reaching near-converged metrics. All estimates are based on a single node with eight H100 GPUs, except for FLUX 1024 Res, which used eight B200 GPUs. The near-converged points are inferred from Figure 4. Estimated training times (in hours) are computed as the number of images iterated (in millions) multiplied by the time-per-million-images values reported in Tables 4 and 5.
Model k imgs to checkpoint k imgs to near convergence Hours to checkpoint Hours to near convergence
SANA 0.6B (512 res) 665 100 3.325 0.5
SANA 1.6B (512 res) 1996 400 9.98 2
SANA 0.6B (1024 res, TrigFlow) 1587 200 22.218 2.8
SANA 1.6B (1024 res, TrigFlow) 1638 150 49.14 4.5
SD3 Medium 1669 200 55.077 6.6
SD3.5 Medium 1269 400 71.064 22.4
SD3.5 Large 696 130 106.488 19.89
Flux 512 Res 778 778 140.818 140.818
Flux 1024 Res 699 300 139.8 60
Appendix G Additional Qualitative Examples
Figure 5: Qualitative results produced by the four-step SiD-DiT generator distilled from FLUX-1.DEV.
@@ Appendix G Additional Qualitative Examples
[…]
Appendix H Prompt Details